# Delayed Scaling of Multi-Type Cosmic F- and D-strings in VOS Models

### v1

@@ 4.4 Physical origin of the ploronged scaling time
This explains why an extremely prolonged scaling time appears in a narrow region of parameter space, precisely where the above cancellation among the coefficients occurs.
From the discussion above, one can identify two essential ingredients in the VOS model that lead to a significantly delayed onset of the scaling regime. The first is the presence of multiple interacting string species (three types in our case), which introduce several terms in the VOS equations with different dependences on the string densities. The second is the existence of free parameters associated with the reconnection probabilities, which allows for the possibility of fine-tuning among these terms. The combination of these two features enables the cancellations responsible for the delayed scaling behavior. This phenomenon can therefore be regarded as a characteristic feature intrinsic to the string network considered in this work. Since these features are also present in the conventional VOS model, they likewise lead to a significantly delayed onset of the scaling regime, as we explicitly confirm in Appendix B.
@@ A.2.1 $P_{3}\ll P_{\text{loop}}$
$\gamma^{2}_{3}=\frac{P_{\text{loop}}^{2}}{2P_{3}}\left\{-\left(1-\frac{2P_{1}}{P_{\text{loop}}}\right)\pm\sqrt{\left(1-\frac{2P_{1}}{P_{\text{loop}}}\right)^{2}+\frac{4P_{3}}{P_{\text{loop}}}}\right\}.$ (81)
@@
$\gamma_{3}^{2}$ $\approx\frac{P_{\text{loop}}^{2}}{2P_{3}}\left\{\left(\frac{2P_{1}}{P_{\text{loop}}}-1\right)+\left(\frac{2P_{1}}{P_{\text{loop}}}-1\right)\right\}$
$=\frac{P_{\text{loop}}^{2}}{P_{3}}\left(\frac{2P_{1}}{P_{\text{loop}}}-1\right)$
$\approx 2\frac{P_{\text{loop}}P_{1}}{P_{3}}\,.$ (82)
@@ A.2.2 $P_{\text{loop}}\ll P_{3}$
$\gamma_{3}^{2}=\frac{P_{3}}{2}\left\{-\left(1-\frac{2P_{1}}{P_{3}}\right)\pm\sqrt{\left(1-\frac{2P_{1}}{P_{3}}\right)^{2}+\frac{4P_{\text{loop}}}{P_{3}}}\right\}\,.$ (85)
@@
$\gamma_{3}^{2}$ $=\frac{P_{3}}{2}\left\{\left(\frac{2P_{1}}{P_{3}}-1\right)+\sqrt{\left(\frac{2P_{1}}{P_{3}}-1\right)^{2}+\frac{4P_{\text{loop}}}{P_{3}}}\right\}$
$\approx P_{3}\left(\frac{2P_{1}}{P_{3}}-1\right)$
$\approx P_{1}\,.$ (86)
@@ C.2.1 $\tilde{P}_{3}\ll\tilde{P}_{\text{loop}}$
$\frac{\tilde{P}_{3}}{2\tilde{P}_{\text{loop}}^{3}}\gamma_{3}^{4}+\frac{\tilde{P}_{3}}{2\tilde{P}_{\text{loop}}^{2}}+\left(1-\frac{2\tilde{P}_{1}}{\tilde{P}_{\text{loop}}}\right)\gamma_{3}^{2}-\tilde{P}_{\text{loop}}^{2}=0.$ (106)
The solution to this quartic equation is
$\gamma_{3}$ $=-\frac{b}{4a}+\frac{1}{2}\sqrt{\frac{b^{2}}{4a^{2}}-\frac{2c}{3a}+\frac{\sqrt[3]{2}(c^{2}-12ad)}{3a(D_{1}+D_{2})^{1/3}}+\frac{(D_{1}+D_{2})^{1/3}}{3\sqrt[3]{2}a}}$
[…]
$~~~~~~~~~~~~~~~\left.+\frac{-\frac{b^{3}}{a^{3}}+\frac{4bc}{a^{2}}}{4\sqrt{\frac{b^{2}}{4a^{2}}-\frac{2c}{3a}+\frac{\sqrt[3]{2}(c^{2}-12ad)}{3a(D_{1}+D_{2})^{1/3}}+\frac{(D_{1}+D_{2})^{1/3}}{3\sqrt[3]{2}a}}}\right)^{1/2}\,$, (107)
where we define
$\left\{~~\begin{array}[]{l}\displaystyle a\equiv\frac{\tilde{P}_{3}}{2\tilde{P}_{\text{loop}}^{3}},\quad b\equiv\frac{\tilde{P}_{3}}{2\tilde{P}_{\text{loop}}^{2}}\\[12.91663pt] \displaystyle c\equiv 1-\frac{2\tilde{P}_{1}}{\tilde{P}_{\text{loop}}},\quad d\equiv\tilde{P}_{\text{loop}}^{2}\\[12.91663pt] D_{1}\equiv 2c^{3}-27b^{2}d+72acd\\[8.61108pt] D_{2}\equiv\sqrt{-4(c^{2}-12ad)^{3}+D_{1}^{2}}\,.\end{array}\right$. (108)
If $\tilde{P}_{3}\ll\tilde{P}_{\text{loop}}\ll\tilde{P}_{1}$, we can approximate
$\frac{c}{a}$ $\approx-2\frac{\tilde{P}_{1}\tilde{P}_{\text{loop}}^{2}}{\tilde{P}_{3}}\,$, (109)
$D_{1}$ $\approx-\frac{2\tilde{P}_{1}}{\tilde{P}_{\text{loop}}}\,$, (110)
$D_{2}$ $\approx 2i\left(\frac{\tilde{P}_{1}}{\tilde{P}_{\text{loop}}}\right)^{3}.$ (111)
Substituting these expressions into (107), we obtain the approximate solution
$\gamma_{3}\approx$ $-\frac{1}{4}\tilde{P}_{\text{loop}}+\frac{1}{4}\tilde{P}_{\text{loop}}\left(\frac{\tilde{P}_{1}}{\tilde{P}_{3}}\right)^{1/2}\left(\frac{32}{3}+\frac{\tilde{P}_{3}}{\tilde{P}_{1}}\right)^{1/2}$
$~~~~~~~+\frac{1}{\sqrt{3}}\tilde{P}_{\text{loop}}\left(\frac{\tilde{P}_{1}}{\tilde{P}_{3}}\right)^{1/2}\left(1+\frac{9}{4}\left(\frac{\tilde{P}_{3}}{\tilde{P}_{1}}\right)^{1/2}+\frac{3}{8}\frac{\tilde{P}_{3}}{\tilde{P}_{1}}\right)$
$\propto$ $~\tilde{P}_{1}^{1/2}\tilde{P}_{3}^{-1/2}\tilde{P}_{\text{loop}}\,.$ (112)
If $\tilde{P}_{1}\ll\tilde{P}_{3}\ll\tilde{P}_{\text{loop}}$ or $\tilde{P}_{3}\ll\tilde{P}_{1}\ll\tilde{P}_{\text{loop}}$, we can approximate
$D_{1}$ $\approx 2\,$, (113)
$D_{2}$ $\approx 0\,.$ (114)
We then obtain
$\gamma_{3}$ $\approx-\frac{1}{4}\tilde{P}_{\text{loop}}+\frac{1}{2}\sqrt{\frac{\tilde{P}_{\text{loop}}^{2}}{4}}+\frac{1}{2}\sqrt{\frac{\tilde{P}_{\text{loop}}^{2}}{2}\left(1-\frac{8\tilde{P}_{\text{loop}}}{\tilde{P}_{3}}\right)+\frac{\tilde{P}_{\text{loop}}}{2}\left(\frac{1}{\tilde{P}_{3}}-1\right)}$
$\propto\tilde{P}_{3}^{-1/2}\tilde{P}_{\text{loop}}.$ (115)
@@ C.2.2 $\tilde{P}_{\text{loop}}\ll\tilde{P}_{3}$
In this case, we can approximate $\gamma_{1}\approx\tilde{P}_{3}$ from (103). Proceeding as in the previous subsection, we obtain the following cubic equation for the scaling value of $\gamma_{3}$ from (104):
$\frac{1}{2\tilde{P}_{3}^{2}}\gamma_{3}^{4}+\frac{1}{2\tilde{P}_{3}}\gamma_{3}^{3}+\left(1-\frac{2\tilde{P}_{1}}{\tilde{P}_{3}}\right)\gamma_{3}^{2}+(\tilde{P}_{3}-\tilde{P}_{\text{loop}})\gamma_{3}-\tilde{P}_{\text{loop}}\tilde{P}_{3}=0.$ (116)
Following the same procedure as for the quartic case, the solution can be expressed as
$\gamma_{3}=$ $-\frac{b}{4a}-\frac{1}{2}\sqrt{\frac{b^{2}}{4a^{2}}-\frac{2c}{3a}+\frac{\sqrt[3]{2}D_{3}}{3a\left(D_{4}+\sqrt{-4D_{3}^{3}+D_{4}^{2}}\right)^{1/3}}+\frac{\left(D_{4}+\sqrt{-4D_{3}^{3}+D_{4}^{2}}\right)^{1/3}}{3\sqrt[3]{2}a}}$
$+\frac{1}{2}\left(\frac{b^{2}}{2a^{2}}-\frac{4c}{3a}-\frac{\sqrt[3]{2}D_{3}}{3a\left(D_{4}+\sqrt{-4D_{3}^{3}+D_{4}^{2}}\right)^{1/3}}-\frac{\left(D_{4}+\sqrt{-4D_{3}^{3}+D_{4}^{2}}\right)^{1/3}}{3\sqrt[3]{2}a}\right$.
$~~~~~~~~~~~~~\left.-\frac{-\frac{b^{3}}{a^{3}}+\frac{4bc}{a^{2}}-\frac{8d}{a}}{4\sqrt{\frac{b^{2}}{4a^{2}}-\frac{2c}{3a}+\frac{\sqrt[3]{2}D_{3}}{3a\left(D_{4}+\sqrt{-4D_{3}^{3}+D_{4}^{2}}\right)^{1/3}}+\frac{\left(D_{4}+\sqrt{-4D_{3}^{3}+D_{4}^{2}}\right)^{1/3}}{3\sqrt[3]{2}a}}}\right)^{1/2}\,$, (117)
where we define
[…]
If $\tilde{P}_{\text{loop}}\ll\tilde{P}_{3}\ll\tilde{P}_{1}$, we can use the approximations
$\frac{c}{a}$ $\approx-2\tilde{P}_{1}\tilde{P}_{3}\,$, (119)
$\frac{d}{a}$ $\approx 2\tilde{P}_{3}^{3}\,$, (120)
$D_{3}$ $\approx\left(\frac{2\tilde{P}_{1}}{\tilde{P}_{3}}\right)^{2}-\frac{3}{2}\,$, (121)
$D_{4}$ $\approx-16\left(\frac{\tilde{P}_{1}}{\tilde{P}_{3}}\right)^{3}+\frac{27}{2}\,.$ (122)
Substituting these expressions yields the approximate solution
$\gamma_{3}\approx$ $-\frac{1}{4}\tilde{P}_{3}-\frac{1}{2}\tilde{P}_{3}+\frac{1}{2}\left(\frac{4}{3}\tilde{P}_{1}\tilde{P}_{3}+4\tilde{P}_{1}\tilde{P}_{3}\right)^{1/2}$
$=$ $\frac{2}{\sqrt{3}}\tilde{P}_{1}^{1/2}\tilde{P}_{3}^{1/2}\left(1-\frac{3\sqrt{3}}{8}\sqrt{\frac{\tilde{P}_{3}}{\tilde{P}_{1}}}\right)$
$\propto$ $~\tilde{P}_{1}^{1/2}\tilde{P}_{3}^{1/2}.$ (123)
If $\tilde{P}_{\text{loop}}\ll\tilde{P}_{1}\ll\tilde{P}_{3}$ or $\tilde{P}_{1}\ll\tilde{P}_{\text{loop}}\ll\tilde{P}_{3}$, we solve (116) perturbatively. The dominant part of the equation becomes
$x^{4}+x^{3}+2x^{2}+2x=0\,$, (124)
where we define $x\equiv\gamma_{3}/\tilde{P}_{3}$. The only real solution is $x_{0}=0$. We therefore expand around $x=x_{0}=0$ as $x=x_{0}+\delta$ and keep terms up to $\mathcal{O}(\delta)$. Substituting this into (116), we obtain
$\displaystyle\delta=\frac{2(1+x_{0})\frac{\tilde{P}_{\text{loop}}}{\tilde{P}_{3}}+4x_{0}^{2}\frac{\tilde{P}_{1}}{\tilde{P}_{3}}}{4x_{0}^{3}+3x_{0}^{2}+4x_{0}+2}.$ (125)
Since $x_{0}=0$, we finally obtain
$\gamma_{3}$ $\approx\tilde{P}_{3}(x_{0}+\delta)=2\tilde{P}_{\text{loop}}$
$\propto\tilde{P}_{\text{loop}}\,.$ (126)
Thus, the asymptotic behavior of $\gamma_{3}$ is

### v2

@@ 4.4 Physical origin of the ploronged scaling time
This explains why an extremely prolonged scaling time appears in a narrow region of parameter space, precisely where the above cancelation among the coefficients occurs.
From the discussion above, one can identify two essential ingredients in the VOS model that lead to a significantly delayed onset of the scaling regime. The first is the presence of multiple interacting string species (three types in our case), which introduce several terms in the VOS equations with different dependences on the string densities. The second is the existence of free parameters associated with the reconnection probabilities, which allows for the possibility of fine-tuning among these terms. The combination of these two features enables the cancelations responsible for the delayed scaling behavior. This phenomenon can therefore be regarded as a characteristic feature intrinsic to the string network considered in this work. Since these features are also present in the conventional VOS model, they likewise lead to a significantly delayed onset of the scaling regime, as we explicitly confirm in Appendix B.
@@ A.2.1 $P_{3}\ll P_{\text{loop}}$
$\gamma^{2}_{3}=\frac{P_{\text{loop}}^{2}}{2P_{3}}\left\{-\left(1-\frac{2P_{1}}{P_{\text{loop}}}\right)+\sqrt{\left(1-\frac{2P_{1}}{P_{\text{loop}}}\right)^{2}+\frac{4P_{3}}{P_{\text{loop}}}}\right\}\,$, (81)
where a negative root is discarded because $\gamma_{3}^{2}$ must be positive.
@@
$\gamma_{3}^{2}$ $\approx 2\frac{P_{\text{loop}}P_{1}}{P_{3}}\,.$ (82)
@@ A.2.2 $P_{\text{loop}}\ll P_{3}$
$\gamma_{3}^{2}=\frac{P_{3}}{2}\left\{-\left(1-\frac{2P_{1}}{P_{3}}\right)+\sqrt{\left(1-\frac{2P_{1}}{P_{3}}\right)^{2}+\frac{4P_{\text{loop}}}{P_{3}}}\right\}\,$, (85)
where a negative root is discarded because $\gamma_{3}^{2}$ must be positive.
@@
$\gamma_{3}^{2}$ $\approx 2P_{1}\,.$ (86)
@@ C.2.1 $\tilde{P}_{3}\ll\tilde{P}_{\text{loop}}$
$\frac{\tilde{P}_{3}}{2\tilde{P}_{\text{loop}}^{3}}\gamma_{3}^{4}+\left(1-\frac{2\tilde{P}_{1}}{\tilde{P}_{\text{loop}}}\right)\gamma_{3}^{2}-\tilde{P}_{\text{loop}}^{2}\approx 0.$ (106)
where we use $\tilde{P}_{3}\ll\tilde{P}_{\text{loop}}$.
If $\tilde{P}_{3}\ll\tilde{P}_{\text{loop}}\ll\tilde{P}_{1}$, the equation can be approximated as
$\frac{\tilde{P}_{3}}{2\tilde{P}_{\text{loop}}^{3}}\gamma_{3}^{4}-\frac{2\tilde{P}_{1}}{\tilde{P}_{\rm loop}}\gamma_{3}^{2}-\tilde{P}_{\text{loop}}^{2}\approx 0\,.$ (107)
Solving this equation yields
$\gamma_{3}\sim 2\tilde{P}_{1}^{1/2}\tilde{P}_{3}^{-1/2}\tilde{P}_{\text{loop}}\propto\tilde{P}_{1}^{1/2}\tilde{P}_{3}^{-1/2}\tilde{P}_{\text{loop}}\,.$ (108)
where the third term in Eq. (107) has been neglected as subdominant.
If $\tilde{P}_{1},\tilde{P}_{3}\ll\tilde{P}_{\text{loop}}$, we can approximate
$\frac{\tilde{P}_{3}}{2\tilde{P}_{\text{loop}}^{3}}\gamma_{3}^{4}+\gamma_{3}^{2}-\tilde{P}_{\text{loop}}^{2}\approx 0\,.$ (109)
Solving this equation yields
$\gamma_{3}\approx\tilde{P}_{\text{loop}}\,.$ (110)
where the first term in Eq. (109) has been neglected as subdominant.
@@ C.2.2 $\tilde{P}_{\text{loop}}\ll\tilde{P}_{3}$
In this case, we can approximate $\gamma_{1}\approx\tilde{P}_{3}$ from (103). Then we obtain the following equation for the scaling value of $\gamma_{3}$ from (104):
$\frac{1}{2\tilde{P}_{3}^{2}}\gamma_{3}^{4}+\frac{1}{2\tilde{P}_{3}}\gamma_{3}^{3}+\left(1-\frac{2\tilde{P}_{1}}{\tilde{P}_{3}}\right)\gamma_{3}^{2}+\tilde{P}_{3}\gamma_{3}-\tilde{P}_{\text{loop}}\tilde{P}_{3}\approx 0.$ (111)
For later convenience, it is useful to rewrite this equation as
$\left[\frac{1}{2}x^{4}+\frac{1}{2}x^{3}+x^{2}+x\right]-\frac{2\tilde{P}_{1}}{\tilde{P}_{3}}x^{2}-\frac{\tilde{P}_{\text{loop}}}{\tilde{P}_{3}}=0.$ (112)
where we have defined $x\equiv\gamma_{3}/\tilde{P}_{3}$.
If $\tilde{P}_{\text{loop}}\ll\tilde{P}_{3}\ll\tilde{P}_{1}$, we first approximate the equation as
$\frac{1}{2}x^{4}-\frac{2\tilde{P}_{1}}{\tilde{P}_{3}}x^{2}\approx 0.$ (113)
Solving this equation yields
$\gamma_{3}\approx 2\tilde{P}_{1}^{1/2}\tilde{P}_{3}^{1/2}\propto$ $~\tilde{P}_{1}^{1/2}\tilde{P}_{3}^{1/2}\,$, (114)
which self-consistently justifies the approximation in (113) $\tilde{P}_{\text{loop}}\ll\tilde{P}_{3}\ll\tilde{P}_{1}$.
If $\tilde{P}_{\text{loop}}\ll\tilde{P}_{1}\ll\tilde{P}_{3}$ or $\tilde{P}_{1}\ll\tilde{P}_{\text{loop}}\ll\tilde{P}_{3}$, the equation reduces to
$\left[\frac{1}{2}x^{4}+\frac{1}{2}x^{3}+x^{2}+x\right]-\frac{\tilde{P}_{\text{loop}}}{\tilde{P}_{3}}=0.$ (115)
Assuming $x\ll 1$, this equation admits the approximate solution
$x\approx\frac{\tilde{P}_{\text{loop}}}{\tilde{P}_{3}},$ (116)
which self-consistently justifies the assumption $x\ll 1$. We therefore obtain
$\gamma_{3}$ $\propto\tilde{P}_{\text{loop}}\,$, (117)
in this regime.
In summary, the asymptotic behavior of $\gamma_{3}$ is